OpenStudy (anonymous):

A paint can has a diameter of 17 centimeters and a height of 24 centimeters. How many cubic centimeters of paint will fill the can?

OpenStudy (across):

You need to know the equation for the volume of a cylinder.

OpenStudy (across):

\[V_{\text{cylinder}}=\pi r^2h,\]where \(r\) is the radius and \(h\) is the height.

OpenStudy (anonymous):

\[V = \pi r ^{2}h\]\[V = 3.14 \times 8.5 cm \times 8.5cm \times 24 cm\]\[V = 5444.76 cm ^{3}\]